# DiffMesh: A Motion-aware Diffusion-like Framework for Human Mesh Recovery from Videos

### v3

@@ C.1 Performance on MPI-INF-3DHP dataset
To conduct experiments on the MPI-INF-3DHP dataset, we follow the same setting as VIBE [20], TCMR [6], and MPS-Net [50], the number of input frames is set to be 16. The input features of each frame are extracted from ResNet-50 [13] without fine-tuning for fair comparisons. The results are shown in Fig. 6. Our DiffMesh consistently outperforms previous methods with significant improvement (more than 5.9 mm $\downarrow$ of MPJPE, 1.8 $\downarrow$ of PA-MPJPE, and 0.7 $\downarrow$ of ACC-ERR). This showcases the remarkable performance enhancement achieved by our approach, highlighting its potential as a state-of-the-art solution for video-based human mesh recovery across various datasets and real-world applications.
@@
C.2 Effectiveness of the number of input frames and the number of additional steps
@@
C.3 Different initial distributions in diffusion
In Section 3.3 of the main paper, we assume that human motion will eventually reach a static state $x_{end}$, which is an initial distribution. In our DiffMesh, we select the template mesh acquired by averaging ground-truth meshes on Human3.6M dataset [30] as the initial distribution. Here we compared the results when using pure Gaussian noise as the initial distribution. The results indicate that selecting template mesh as the initial distribution is a better choice than Pure Gaussian noise for the video-based HMR task.
Table 8: Evaluation of different initial distributions on the 3DPW dataset.
3DPW
[…]
Pure Gaussian noise 87.0 75.8 46.7 6.1
Template mesh 86.4 75.7 45.6 6.1
C.4 Different design choices of our transformer-based diffusion model
@@
Table 9: Ablation study of transformer block design on 3DPW dataset.
@@
C.5 Effectiveness of the auxiliary loss:
@@
$\mathcal{L}_{HMR}$ 86.8 76.0 47.1 6.2
$\mathcal{L}_{HMR}+\mathcal{L}_{aux}$ 86.4 75.7 45.6 6.1
@@
Appendix D Human Mesh Visualization
We first visualize the qualitative comparison on the 3DPW [48] dataset in Fig. 10. The circle areas highlight locations where our DiffMesh performs better than SOTA GLoT [42].

### v4

@@ C.1 Performance on MPI-INF-3DHP dataset
To conduct experiments on the MPI-INF-3DHP [36] dataset, we follow the same setting as VIBE [22], TCMR [5], and MPS-Net [54] . The input features of each frame are extracted from ResNet-50 [14] without fine-tuning for fair comparisons. The results are shown in Fig. 5. Our DiffMesh consistently outperforms previous methods with significant improvement (more than 5.9 mm $\downarrow$ of MPJPE, 1.8 $\downarrow$ of PA-MPJPE, and 0.7 $\downarrow$ of ACC-ERR). This showcases the remarkable performance enhancement achieved by our approach, highlighting its potential as a state-of-the-art solution for video-based human mesh recovery across various datasets and real-world applications.
C.2 Effectiveness of the number of input frames and additional steps
@@
C.3 Different design choices of our transformer-based diffusion model
@@
Table 7: Ablation study of transformer block design on 3DPW dataset.
C.4 Effectiveness of the auxiliary loss:
@@
Without $\mathcal{L}_{aux}$ 86.8 76.0 47.1 6.2
With $\mathcal{L}_{aux}$ 86.4 75.7 45.6 6.1
C.5 Inference Time Analysis:
Methods like MPS-Net [54] and GLoT [45] only estimate the human mesh of the center frame given 16 frames as their input. Considering these methods can extract all features by their backbone once and then utilize batch processing to accelerate the inference speed, we provide a more thorough inference time comparison in Table 9.
In this experiment, the video input comprises a total of 64 frames. Upon feature extraction from the backbone (with the shape of $[64,2048]$), MPS-Net and GLoT require the creation of 64 batch input tubes $[64,16,2048]$ through padding and sliding window. Since their models only return the output mesh of the center frame, the output would be $[64,1,6890,3]$, indicating output mesh vertices $[6890,3]$ across 64 frames. In contrast, our DiffMesh just needs to reshape the input $[64,2048]$ into 4 batches, resulting in the shape of $[4,16,2048]$. Consequently, the output of DiffMesh is $[4,16,6890,3]$, which is then reshaped back to $[64,6890,3]$. Based on the total processing time, our DiffMesh is more efficient than MPS-Net [54] and GLoT [45] since DiffMesh can output human meshes of all input frames.
Video-based
Methods
total frames for
video input
Backbone
features after backbone are
reshaped for model processing
Output
shape
processing time
(without backbone time)
MPS-Net [54] 64 ResNet50 [14] [64,2048] to [64,16,2048] [64,1,6890,3] to [64,6890,3] 1.04 s
GLoT [45] 64 ResNet50 [14] [64,2048] to [64,16,2048] [64,1,6890,3] to [64,6890,3] 1.17 s
DiffMesh (ours) 64 ResNet50 [14] [64,2048] to [4,16,2048] [4,16,6890,3] to [64,6890,3] 0.34 s
Table 9: Inference time comparison on 3DPW dataset between our DiffMesh and previous video-based HMR methods with the same hardware platform ( single NVIDIA A5000 GPU is used).
Appendix D Human Mesh Visualization
We first visualize the qualitative comparison on the 3DPW [51] dataset in Fig. 11. The circle areas highlight locations where our DiffMesh performs better than GLoT [45].
In our experimental setup, we utilize 16 input frames, and the total number of steps is set to 30. In the reverse motion process, DiffMesh outputs [$y_{1},y_{2}\cdots,y_{30}$] over 30 steps. For the output mesh sequence of 16 frames, we use [$y_{1},y_{2}\cdots,y_{16}$]. Additionally, we generate the mesh from [$y_{16},y_{17}\cdots,y_{30}$], as visually depicted in Fig 13. This visualization illustrates the trend of the generated human mesh gradually decoding toward the desired human mesh of each input frame.
Appendix E Broader impact and limitation
DiffMesh establishes an innovative connection between diffusion models and human motion, facilitating the generation of accurate and temporal smoothness output mesh sequences by integrating human motion into both the forward and reverse processes of the diffusion model. By enabling direct 3D human mesh reconstruction from 2D video sequences, DiffMesh eliminates the dependency on additional motion sensors and equipment, thereby streamlining the process and reducing costs.
However, despite its advancements, DiffMesh is not without limitations. Similar to previous methods, DiffMesh may face challenges in scenarios with substantial occlusions, resulting in the production of unrealistic mesh outputs. To address this issue, further exploration into spatial-temporal interactions within the human body is warranted, serving as a focal point for our future research. Additionally, DiffMesh may encounter difficulties in rare and complex pose scenarios due to the constraints of limited training data, highlighting the necessity for ongoing development and refinement efforts.
@@ Appendix E Broader impact and limitation
Figure 13: Visualization of decoding steps during the reverse motion process.